What is the osmotic pressure of a .333M solution of C9H12O6 at 25 Celsius?
What is the freezing point of a 1.77m solution of CBr4 in C6H6?
What is the boiling point of a 2.50m solution of C6H4Cl2 (non volatile)in CCl4?
A solution of C2H2O4(non volatile) in C2H5OH has a boiling point of 80.38C. What is the morality of solution?
1
Expert's answer
2020-11-23T06:59:28-0500

A) Using osmotic pressure formula ,n=inRT/V​=iMRTFor molarity M=1000+m×Mw1000m×d​putting the values =(1000×1.25×342)
(1000×1.25×1.34)​=1.733MHence,n=iMRTn=28.6atm    (For:i=1)


B) The osmotic pressure of same solution is 0.474 atm
ΔTf​=5.5−5.4=0.1oC
π=CM​ST(C+M=molarity)
ΔTf=Cm×Kf(Cm=molality)for dilute solutions CM​=Cm​
∴π=Kf​
ΔTf​×ST=4.9
0.1×0.0821×283=0.474atm